# WaterPark: A Robustness Assessment of Language Model Watermarking

### v1

@@ Abstract
To mitigate the misuse of large language models (LLMs), such as disinformation, automated phishing, and academic cheating, there is a pressing need for the capability of identifying LLM-generated texts. Watermarking emerges as one promising solution: it plants statistical signals into LLMs’ generative processes and subsequently verifies whether LLMs produce given texts. Various watermarking methods (“watermarkers”) have been proposed; yet, due to the lack of unified evaluation platforms, many critical questions remain under-explored: i) What are the strengths/limitations of various watermarkers, especially their attack robustness? ii) How do various design choices impact their robustness? iii) How to optimally operate watermarkers in adversarial environments?
To fill this gap, we systematize existing LLM watermarkers and watermark removal attacks, mapping out their design spaces. We then develop WaterPark, a unified platform that integrates 10 state-of-the-art watermarkers and 12 representative attacks. More importantly, leveraging WaterPark, we conduct a comprehensive assessment of existing watermarkers, unveiling the impact of various design choices on their attack robustness. For instance, a watermarker’s resilience to increasingly intensive attacks hinges on its context dependency. We further explore the best practices to operate watermarkers in adversarial environments. For instance, using a generic detector alongside a watermark-specific detector improves the security of vulnerable watermarkers. We believe our study sheds light on current LLM watermarking techniques while WaterPark serves as a valuable testbed to facilitate future research. [^1] All the source code and data are publicly available: https://github.com/JACKPURCELL/sok-llm-watermark
@@ 1 Introduction
The recent advances in large language models (LLMs), including GPT[46], Claude[3], Mistral[22], and Llama[58], have significantly enhanced our capabilities of general-purpose text generation and complex problem-solving. However, such unprecedented capabilities also give rise to a variety of possible misuses, such as disinformation[21], automated phishing[4], and academic cheating[7]. There is thus a pressing need for the capability of identifying LLM-generated content.
A simple approach is to train classifiers that rely on the intrinsic difference between LLM- and human-generated texts[42]. Yet, with LLMs becoming increasingly capable, the gap between LLM- and human-generated texts is narrowing. Thus, watermarking is gaining attention as a promising alternative. By embedding specific statistical signals (“watermarks”) into LLMs’ generative processes, it is possible to verify whether given texts are produced by such models, as illustrated in Figure 1. A variety of LLM watermarking methods (“watermarkers") have been proposed[1, 35, 36, 26, 20, 76, 29], each with unique design choices and desirable properties, raising a set of intriguing questions:
Figure 1: LLM watermarking and watermark removal attacks.
RQ1 - What are the strengths/limitations of various watermarkers, especially their robustness against manipulations?
RQ2 - How do different design choices impact the attack robustness of watermarkers?
RQ3 - What are the best practices for operating watermarkers in adversarial environments?
Despite recent benchmarking studies on LLM watermarking, existing research is limited in addressing these questions. For instance, WaterBench[59] primarily focuses on watermarking effectiveness; MarkMyWords[49] mainly evaluates the robustness of a specific watermarker[26]; MarkLLM[47] focuses on providing a platform to compare watermark detectability, basic robustness, and text quality. Moreover, these studies lack an in-depth analysis of how a watermarker’s design choices impact its robustness. Consequently, the aforementioned questions remain largely unexplored due to the absence of unified, comprehensive evaluation platforms.
@@
Our Work. To bridge this critical gap, this work conducts a systematic study of state-of-the-art LLM watermarkers, focusing on their robustness against watermark removal attacks. We aim to understand how various design choices affect watermarkers’ attack resilience and identify best practices for operating watermarkers in adversarial environments. We provide insights through three key aspects:
Platform –- We develop WaterPark, the first open-source platform dedicated to evaluating the attack robustness of LLM watermarkers in a unified and comprehensive manner. As of 09/04/2024, WaterPark integrates 10 state-of-the-art watermarkers, 12 representative watermark removal attacks, and 8 key metrics. Moreover, WaterPark offers a comprehensive suite of tools for in-depth robustness assessment, including next-token distribution comparisons, attack combination analyses, and what-if scenario evaluations
Assessment –- Leveraging WaterPark, we empirically evaluate the attack resilience of representative LLM watermarkers, leading to many interesting findings, many of which challenge the conclusions in prior work, as summarized in Table 1. For instance, i) UPV is more prone to false positives than they claim, making it challenging to achieve an ideal TPR at lower FPR; ii) Some watermarks, such as UPV, may perform poorly in a single paraphrase but can still be detected after multiple paraphrases. iii) SIR claims that they can withstand paraphrase attacks, but they are unable to resist high-intensity dipper attacks or paraphrases generated by GPT. iv) UB, in addition to being resistant to token replacement attacks as mentioned in the original text, demonstrates poor robustness against other types of attacks.
Exploration – We also explore how a watermarker’s design choices impact its attack robustness, unveiling critical trade-offs between different types of robustness. For instance, i) watermarkers that rely on explicit statistical signals tend to be robust against linguistic variation attacks; ii) watermarks based on distribution transform show great resistance to attacks that mix watermarked and non-watermarked texts; iii) watermarkers’ robustness against increasingly intensive attacks largely depends on their context dependencies. Further, we also explore the best practices to operate watermarkers in adversarial environments. For instance, i) employing a generic detector alongside a watermark-specific detector improves the security of vulnerable watermarkers; ii) evaluating watermarkers’ robustness needs to consider not only individual attacks but also their combinations; iii) paraphrasing using highly capable LLMs serves as a dominant attack to test watermarkers’ robustness lower-bounds.
This study sheds light on the current LLM watermarking techniques while WaterPark servers as a useful platform to facilitate future research in this space.
2 Preliminaries
We first introduce fundamental concepts and assumptions used throughout the paper.
2.1 LLM Watermarking
@@
At a high level, a watermark is a pattern embedded in a given signal (e.g., image and text) for identifying the signal’s source. In the context of LLMs, watermarks can be used to prove that a given text is LLM-generated (or even generated by specific LLMs). Typically, an LLM watermarking method (“watermarker”) comprises three components: the LLM, watermarking procedure (generator), and detection procedure (detector), as illustrated in Figure 1.
@@
Typically, the generator produces watermarked texts iteratively. At each iteration, with access to a secret key $k$, the previous tokens $x_{<t}$, and the LLM’s next-token distribution $p(x_{t}|x_{<t})$, the generator generates a perturbed distribution $\tilde{p}(x_{t}|x_{<t})$, from which the next token $x_{t}$ is sampled. Meanwhile, with access to a secret key $k^{\prime}$, the detector determines whether $M$ generates a given text. Note that in symmetric schemes, $k=k^{\prime}$, while in asymmetric schemes[11], $k\neq k^{\prime}$. This study mainly focuses on symmetric watermarking schemes due to their widespread adoption.
2.2 Desiderata of Watermarking
We consider that an ideal LLM watermarker possesses the following desirable properties.
Effectiveness – The detection of (non-)watermarked texts should be highly accurate;
Fidelity – The watermarking should minimally impact the original texts’ semantics;
Robustness – The watermarked texts should be resilient against potential watermark removal attacks.
Besides, existing studies[5] also advocate indistinguishability – the watermarked and non-watermarked texts should be indistinguishable, a quality often measured by the distributional difference between the original texts and the expectation of the watermarked texts. However, recent work[49] regards this property as overly restrictive, prioritizing criteria such as effectiveness and fidelity above indistinguishability. In this study, we primarily focus on examining the robustness of current watermarking techniques and their interplay with other evaluation criteria.
2.3 Watermark Removal Attacks
@@
One critical property of LLM watermarkers is their robustness against potential attacks. This study focuses on watermark removal attacks[25] designed to destroy the watermarks embedded in given watermarked texts while preserving their semantics. We assume a threat model similar to prior work[49, 76, 52, 27], as shown in Figure 1. The adversary intercepts the watermarked text $\tilde{T}$, modifies it, and then forwards the altered text $T^{\prime}$ to the watermark detector, aiming to avoid detection.
Adversary’s goals. The adversary modifies the watermarked text $\tilde{T}$ as $T^{\prime}$ such that the following objectives are met: effectiveness – $T^{\prime}$ evades the watermark detector (i.e., detected as non-watermarked) and quality – $T^{\prime}$ preserves $\tilde{T}$’s original semantics.
Adversary’s knowledge. Following prior work[49, 76, 52, 27], we assume the adversary has access to sample watermarked and non-watermarked texts, but cannot reproduce the watermarking procedure or interact with the detection procedure. Otherwise, it is trivial to launch adversarial attacks against the detector (e.g., by identifying and replacing green/red-listed tokens[26]).
Adversary’s capability. We assume the adversary can i) modify the watermarked text in a computationally efficient way (e.g., synonym substitution), ii) use LLMs less capable than the target LLM, and/or iii) train a detector based on given watermarked and non-watermarked texts. We argue that these assumptions are realistic and practical. Otherwise, using capable LLMs may easily generate highly-quality, non-watermarked texts to evade detection, while launching computationally expensive attacks may significantly increase the adversary’s cost.
Index-dep. Text-dep. Context-free Dist.-shift Dist.-reweight Dist.-transform Score-based Diff.-based Model-based
[…]
Table 2: A taxonomy of LLM watermarkers.
3 Taxonomies
3.1 Context Dependency
$r_{t}=f_{k}(t),$ (1)
that only depends on the index $t$ of the next token; $r_{t}$ is then used to generate the perturbation $\Delta_{t}$[29].
Text-dependent watermark. The watermarker considers the previous tokens $x_{<t}$ as the context window to generate $\Delta_{t}$[26, 1, 20]. For instance, GO[1] computes the hash of the concatenation of previous $w$ tokens as $r_{t}$:
$r_{t}=f_{k}(x_{t-w}\|\ldots\|x_{t-1}),$ (2)
while TGRL[26] suggests using the min hash of the previous token:
$r_{t}=\min(f_{k}(x_{t-w}),\ldots,f_{k}(x_{t-1}))$ (3)
Similarly, UB[20] concatenates the previous $w$ tokens to generate the context code for reweighting the logits. UPV[35] and SIR[36] use neural networks to generate $\Delta_{t}$ based on the previous tokens.
Context-free watermark. The watermarker applies a universal perturbation $\Delta_{t}$ across the next-token distributions of all the tokens without considering their contexts[76].
3.2 Generation Strategy
Distribution shift. TGRL[26] modifies $p(x_{t}|x_{<t})$ by adding a shift $\delta$ to the logits of “green-list” tokens (the remaining as “red-list” tokens) as the modified distribution $\tilde{p}(x_{t}|x_{<t})$. A token $x$ is considered as green listed if $\pi_{r_{t}}(x)<\gamma d$ where $\pi_{r_{t}}$ is a permutation seeded by $r_{t}$, $\gamma$ is a parameter to control the size of the green list, and $d$ is the number of vocabulary size. Similarly, UG[76] uses a fixed red-green split over the vocabulary, showing greater robustness than TGRL against edit-distance-bounded attacks due to its “hard” split. UPV[35] selects the top-$k$ tokens from $p(x_{t}|x_{<t})$, applies a neural network to predict the green-list tokens from these tokens, and adds $\delta$ to the logits of green-list tokens. Unlike the other strategies, the distribution-shift strategy preserves the diversity of generated tokens; however, it can not be made indistinguishable, as $\tilde{p}(x_{t}|x_{<t})$ and $p(x_{t}|x_{<t})$ are inherently distinguishable.
Distribution reweight. Similar to distribution shift, this strategy alters the next-token distribution but uniquely perturbs the logit of each token. For instance, SIR[36] trains a neural network to predict the perturbation to logit of each token, which is unbiased (no preference over specific tokens) and balanced (with total perturbation summing up to 0). UB[20] advocates unbiased watermarking such that the expectation of the reweighted distribution agrees with the original distribution. It proposes two reweighting schemes: $\delta$-reweighting uniformly samples a token from $p(x_{t}|x_{<t})$ and changes its probability to 1; $\gamma$-reweighting shuffles all the tokens, rejects the first half, and double the probabilities of the remaining half. DIP[66] also uses a distribution-reweight generation strategy similar to UB but does not need to access the LM during detection.
Distribution transform. Another line of watermarkers apply randomized transform on $p(x_{t}|x_{<t})$ to sample $x_{t}$. For instance, RDF-EXP[29] and GO[1] use the Gumbel-max trick and apply the exponential transform. Let $p(x_{t}|x_{<t})=\{p_{i}\}_{i=1}^{d}$ be the distribution over the next token $x_{t}$. Then $x_{t}$ is sampled as:
$x_{t}=\arg\max_{i\leq d}r_{i}^{1/{p_{i}}}$ (4)
where $r_{i}$ is generated by the pseudo-random functions in Eq. 1 and Eq. 2. Similarly, RDF-IST[29] applies inverse transform over $p(x_{t}|x_{<t})$. With $\pi_{k}$ as a random permutation seeded by the secret key $k$, the next token $x_{t}$ is selected as:
3.3 Detection method
Score-based detection. The detector computes the random value $r_{i}$ for each position, the per-token statistics $s(x_{i},r_{i})$, and a score over such statistics, which is then subjected to a one-tailed statistical test to determine whether the text is watermarked. The per-token statistics vary with the concrete generators. For instance, GO[1] defines $s(x_{i},r_{i})=-\log(1-h_{r_{i}}(x_{i}))$, while TGRL[26] defines $s(x_{i},r_{i})=1$ if $x_{i}$ is in the green list and 0 otherwise. One simple way to aggregate the per-token statistics is to compute their sum[1, 26, 76, 36]:
$S=\sum_{i=1}^{n}s(x_{i},r_{i})$ (6)
However, as the random values may be misaligned with the tokens (e.g., due to editing), a more robust way is to compute the alignment score (e.g., edit score[29, 60]):
$S=s^{\psi}(n,n)\;\text{s.t.}\;s^{\psi}(i,j)=\min\left\{\begin{array}[]{l}s^{\psi}(i-1,j-1)+s(x_{i},r_{j})\\ s^{\psi}(i,j-1)+\psi\\ s^{\psi}(i-1,j)+\psi\end{array}\right.$ (7)
where $\psi$ is the “edit-cost” parameter.
Differential-based detection. This line of detectors also relies on the score of a given text. However, the score is computed by comparing the given text with the non-watermarked text generated by the same LLM. For instance, UB[20] computes the log-likelihood ratio (LLR) score:
$s(i)=\log\frac{\tilde{p}(x_{i}|x_{<i})}{p(x_{i}|x_{<i})}$ (8)
Model-assisted detection. Instead of computing the per-token statistics, which are often subject to watermark removal attacks, one may also train a model to predict whether the given text is watermarked. UPV[35] trains a neural network, which shares the same embedding layers with its generator, to detect watermarked texts. Similarly, one may develop a generic detector by training it to distinguish watermarked and non-watermarked texts. We explore this option in §5.
3.4 Watermark removal attacks
@@
Linguistic variation attack. This class of attacks perturb the linguistic features of the watermarked text, without changing its semantics. We consider the set of perturbations in HELM[34], which simulate natural linguistic variations encountered in human interactions with text typing interfaces: i) Lowercasing converts all the words to lower-cases, which potentially affects the interpretation of proper nouns or emphases. ii) Contracting/expanding replaces phrases with their contracted/expanded forms (e.g., “I am” to “I’m” and vice versa), which may impact the tokenizer.
Lexical editing attack. This class of attacks modifies individual words, aiming to maintain the original text’s semantics. Specifically, we consider the following editing operations: i) Misspelling, similar to text-bugger[32], replaces words with their common misspellings (plural forms also considered); ii) Typoing replaces certain letters in a word with others; iii) Synonymizing replaces words with their synonyms using WordNet[43]; and iv) Swapping randomly exchanges the positions of two words within the text, which alters the text structure while potentially preserving the overall semantics.
Text-mixing attack. This class of attacks aims to “dilute” the watermark by mixing the watermarked text with non-watermarked text fragments. Specifically, the copy-pasting attack[27] embeds the watermarked text into the context of non-watermarked, human-written text. Note that the influence of non-watermarked text can be controlled by setting the fractions of watermarked and non-watermarked text fragments (i.e., the mixing weights).
Paraphrasing attack. This class of attacks relies on an additional LLM (i.e., paraphraser) to re-write the given watermarked text to evade the detector. For instance, a light paraphraser[52] (e.g., T5-based paraphraser[8]) can paraphrase the watermarked text sentence-by-sentence, while a more capable paraphraser (e.g., DIPPER[28]) paraphrases the watermarked text in one-shot, also enabling to control lexical diversity and token order diversity.
Similar to paraphrasing, the translating attack uses a translator LLM (e.g., Seamless [6]) to cycle the watermarked text through multiple languages (e.g., from English to French and back to English). This process can significantly alter the sentence structure and phrasing.
Black-box adversarial attack. In this class of attacks, we assume the adversary has access to a generic detector that is trained to distinguish watermarked and non-watermarked texts. The adversary then perturbs the watermarked text based on this surrogate detector. We explore this attack type in §6.
4 Platform
We implement representative watermarkers and watermark removal attacks on WaterPark, a platform for evaluating different watermarkers in a comprehensive and unified manner. With a modular design, WaterPark can be easily extended to incorporate new watermarkers and attacks.
4.1 Watermarkers and Attacks
Currently, WaterPark incorporates all the watermarkers in Table 2, including 8 one-bit and 2 multi-bit watermarkers. By default, the generator and detector of each watermarker are configured following its official documentation. In addition to the watermark-specific detector, we also implement a generic detector. Specifically, we approach the task as a binary classification problem. We gather watermarked and non-watermarked texts as positive and negative samples. The target LLM and the watermark generator jointly generate the positive samples, while the negative samples are non-watermarked texts generated by ChatGPT-3.5-Turbo (under the same prompt). We train a RoBERTa-based[39] binary classifier as the detector using the training data.
Meanwhile, we implement all the watermark removal attacks in Table 3. Specifically, the linguistic variation attacks follow the definitions of HELM[34]; the text-mixing attack is instantiated as the copy-pasting[27] attack; and the paraphrasing attack is instantiated as the Dipper[28] attack.
4.2 LLMs and Datasets
To factor out the influence of individual LLMs and datasets, WaterPark supports a range of LLMs and datasets.
Specifically, WaterPark integrates a variety of LLMs including OPT-1.3B[74], Llama2-7B[58], Mistral-7B[22], and Gemma-2B/7B[15]; it also includes a variety of datasets including: C4 is a cleaned version of Common Crawl’s web crawl corpus, which is used to train the T5 text-to-text Transformer models; HC3 (Human ChatGPT Comparison Corpus) consists of around 40K questions and their corresponding human/ChatGPT answers; LFQA (Long-Form QA)[13] is a long-form question-answering dataset that consists of around 24K question and answer pairs.
4.3 Metrics
WaterPark evaluates a given watermarker from multiple complementary aspects.
Fidelity. To evaluate the impact of watermarking on text quality, WaterPark employs the following metrics to measure the difference between the original text $T$ and the watermarked text $\tilde{T}$.
BLEU[48] measures the lexical similarity of $T$ and $\tilde{T}$ by calculating the proportion of $n$-grams matched in $\tilde{T}$ and $T$.
BERTScore[75] measures the token-level similarity of $T$ and $\tilde{T}$ by leveraging the pre-trained contextual embeddings from BERT[10] and matching tokens in $T$ and $\tilde{T}$ using cosine similarity.
P-SP[65] evaluates the semantic similarity of $T$ and $\tilde{T}$ using the cosine similarity of their encodings. In particular, the encoding of $T$ (or $\tilde{T}$) is calculated by averaging the embeddings of its subword units generated by SentencePiece[30].
MAUVE[50] compares the distributions of $T$ and $\tilde{T}$ by computing an information divergence curve within a quantized embedding space. The area under this divergence curve provides a scalar summary of the trade-off between Type I ($T$ places high mass in areas where $\tilde{T}$ has low mass) and Type II (vice versa) errors.
5 Evaluation
Leveraging WaterPark, we empirically assess representative LLM watermarkers, focusing on attack robustness and its relationships to other criteria. To our knowledge, this represents the most comprehensive evaluation of state-of-the-art watermarkers to date.
5.1 Experimental Setting
We implement and configure all watermarkers and attacks in accordance with their official documentation. Table 6 summarizes the default parameter settings. To validate our implementation, we assess their effectiveness and fidelity, as reported in §B and §C, corroborating results from existing literature. All experiments are conducted on a workstation running Nvidia RTX A6000 GPUs and 256GB of memory.

### v2

@@ Abstract
Various watermarking methods (“watermarkers”) have been proposed to identify LLM-generated texts; yet, due to the lack of unified evaluation platforms, many critical questions remain under-explored: i) What are the strengths/limitations of various watermarkers, especially their attack robustness? ii) How do various design choices impact their robustness? iii) How to optimally operate watermarkers in adversarial environments? To fill this gap, we systematize existing LLM watermarkers and watermark removal attacks, mapping out their design spaces. We then develop WaterPark, a unified platform that integrates 10 state-of-the-art watermarkers and 12 representative attacks. More importantly, by leveraging WaterPark, we conduct a comprehensive assessment of existing watermarkers, unveiling the impact of various design choices on their attack robustness. We further explore the best practices to operate watermarkers in adversarial environments. We believe our study sheds light on current LLM watermarking techniques while WaterPark serves as a valuable testbed to facilitate future research. [^1] All the source code and data are publicly available: https://github.com/JACKPURCELL/sok-llm-watermark
WaterPark: A Robustness Assessment of Language Model Watermarking
Jiacheng Liang Stony Brook University                        Zian Wang Stony Brook University                        Lauren Hong Stony Brook University
Shouling Ji Zhejiang University                        Ting Wang Stony Brook University
@@ 1 Introduction
The recent advances in large language models (LLMs), including GPT (Openai, ) and Llama (Touvron et al., 2023), have significantly enhanced our capabilities of general-purpose text generation and complex problem-solving, but also raised concerns about misuse through disinformation (Izacard et al., 2022), phishing (Bender et al., 2021), and academic dishonesty (Compilatio, ). There is thus a pressing need for the capability of identifying LLM-generated content.
A simple approach is to train classifiers to distinguish between LLM- and human-generated texts (Lütkebohle, ). However, as LLMs improve, this distinction becomes less clear. Watermarking has emerged as an alternative solution, embedding statistical signals (“watermarks”) during generation to verify LLM-produced texts. Various watermarking methods (“watermarkers") have been developed (Aaronson and Kirchner, ; Liu et al., 2023a, 2024; Kirchenbauer et al., 2023a; Hu et al., 2024; Zhao et al., 2023; Kuditipudi et al., 2023), each with unique design choices and desirable properties, raising a set of intriguing questions:
Figure 1: Illustration of LLM watermarking and watermark removal attacks.
RQ1 – What are the strengths and limitations of various watermarkers, especially their robustness against manipulations?
RQ2 – How do different design choices impact the attack robustness of watermarkers?
RQ3 – What are the best practices for operating watermarkers in adversarial environments?
@@
Despite recent efforts to benchmark LLM watermarkers, existing research is limited in addressing these questions. WaterBench (Tu et al., 2023) primarily focuses on watermarking effectiveness; MarkMyWords (Piet et al., 2023) mainly evaluates the robustness of a specific watermarker (Kirchenbauer et al., 2023a); MarkLLM (Pan et al., 2024) focuses on providing a platform to compare watermark detectability, basic robustness, and text quality. Zhao et al. (2024) provides a comprehensive overview of watermarking techniques. Moreover, these studies lack an in-depth analysis of how a watermarker’s design choices impact its robustness. Consequently, the aforementioned questions remain largely unexplored.
To bridge this gap, this work conducts a systematic study of state-of-the-art LLM watermarkers, focusing on their attack robustness. We aim to understand how various design choices affect attack resilience and identify best practices for operating watermarkers in adversarial environments.
We develop WaterPark, the first open-source platform dedicated to evaluating the attack robustness of LLM watermarkers in a unified and comprehensive manner. As of 12/15/2024, WaterPark integrates 10 state-of-the-art watermarkers, 12 representative watermark removal attacks, and 8 key metrics. Moreover, WaterPark offers a comprehensive suite of tools for in-depth robustness assessment, including next-token distribution comparisons, attack combination analyses, and what-if scenario evaluations. Leveraging WaterPark, we empirically evaluate the attack resilience of representative LLM watermarkers, leading to many interesting findings, which challenge the conclusions in prior work, as summarized in Table 1. We also explore how a watermarker’s design choices impact its attack robustness, unveiling critical trade-offs between different types of robustness.
2 LLM Watermarking
@@
Conceptually, a watermark is a pattern embedded in a given signal (e.g., text) for identifying the signal’s source. In the context of LLMs, watermarks can be used to prove that a given text is LLM-generated (or even generated by specific LLMs). An LLM watermarking method (“watermarker”) comprises three components: the LLM, watermarking procedure (generator), and detection procedure (detector), as shown in Figure 1.
@@
Typically, the generator produces watermarked texts iteratively. Note that in symmetric schemes, $k=k^{\prime}$, while in asymmetric schemes (Fairoze et al., 2023), $k\neq k^{\prime}$. This study mainly focuses on symmetric watermarking schemes due to their widespread adoption.
The current watermarkers can be categorized based on the information carried by the watermarks (e.g., one-bit versus multi-bit) and their key design choices, including context dependency, generation strategy, and detection method. The key design factors of each category are deferred to §B .
3 Platform
3.1 Threat Model
@@
One critical property of LLM watermarkers is their robustness against potential attacks. We assume a threat model similar to prior work (Piet et al., 2023; Zhao et al., 2023; Sankar Sadasivan et al., 2023; Kirchenbauer et al., 2023b), as shown in Figure 1. We assume the adversary has access to sample watermarked and non-watermarked texts, but cannot reproduce the watermarking procedure or interact with the detection procedure. The adversary modifies the watermarked text $\tilde{T}$ as the altered text $T^{\prime}$ such that the following objectives are met: effectiveness – $T^{\prime}$ evades the watermark detector (i.e., detected as non-watermarked) and quality – $T^{\prime}$ preserves $\tilde{T}$’s original semantics.
3.2 Metrics
Fidelity. To evaluate the impact of watermarking on text quality, WaterPark employs the following metrics to measure the difference between the original text $T$ and the watermarked text $\tilde{T}$. We employ WER (word error rate), BLEU (Papineni et al., 2002), BERTScore (Zhang et al., 2020) and P-SP (Wieting et al., 2022) as metrics to assess fidelity. Detailed descriptions of these metrics can be found in §C. Note that these metrics are also used to measure the impact of watermark removal attacks on the quality of the modified text $\tilde{T}^{\prime}$, relative to the watermarked text $\tilde{T}$.
4 Evaluation
We leverage WaterPark to empirically assess representative LLM watermarkers, focusing on their attack robustness and other related criteria.
4.1 Experimental Setting
We implement and configure all watermarkers and attacks in accordance with their official documentation. Table 5 summarizes the default parameter settings. To validate our implementation, we assess their effectiveness and fidelity, as reported in §E and §F, corroborating results from existing literature. A detailed description of each attack can be found in §D. We measure the TPRs of different watermarkers against each attack (with their FPRs fixed as 1%).